1. The lawn sprinkler consists of four arms that rotate in the horizontal plane. The diameter of each nozzle is 8 mm and the water is supplied through the hose at 0.006 m3/s and is ejected through the four nozzles into the atmosphere. Determine the torque required to keep the arms from rotating.



2. Water is discharged into the atmosphere from the pipe at 10 m/s. Determine the horizontal and vertical components of force and the moment that is developed at the fixed support A in order to hold the pipe in equilibrium. Neglect the resistance provided by the pipe at C, and the weight of the pipe and the water within it.


1
Expert's answer
2021-12-16T18:03:36-0500

1.

"Q=\frac{Q_0}4=0.0015~\frac{m^3}s,"

"v=\frac QA=\frac Q{\pi r^2}=29.8~\frac ms,"

"\omega=\frac v l=99.5~\frac {rad}s."

2.

"F=pA,"

"A=\pi r^2,"

"p=\frac{\rho v^2}2,"

"F=\frac{\pi r^2\rho v^2}2=855~N."